OpenStudy (anonymous):

The rectangular bird sanctuary with one side along a straight river is to be constructed so that it contains 8 km^2 of area. Find the dimensions of the rectangle to minimize the amount of fence necessary to enclose the remaining three sides.

OpenStudy (lochana):

you want to apply calculus for this. have you learned calculus before?

OpenStudy (anonymous):

Yes Can you help me derive an equation

OpenStudy (lochana):

yes. You can take the advantage of calculus at this point. fist let's draw our sanctuaryd say it has x and y sides

OpenStudy (lochana):

Area(A) = xy = 8

OpenStudy (lochana):

sum of distance of fence(D) = 2x + y

OpenStudy (lochana):

now lets take D in terms of x using above 2 equations

OpenStudy (lochana):

D = 2x + 8/x

OpenStudy (lochana):

now this is the distance of fence. you need to find minimum of it. so we take first derivative of D with respect to x

OpenStudy (anonymous):

ok

OpenStudy (lochana):

So d(D)/d(x) = 2 - 8/x^2

OpenStudy (lochana):

and you will get \[\frac{ d(D) }{ d(x) } = \frac{ 2(x-2)(x+2) }{ x^2 }\]

OpenStudy (lochana):

now draw the graph. 

OpenStudy (anonymous):

I lost you for how you got your equation for d/x

OpenStudy (anonymous):

how is it 2-(8/x^2)

OpenStudy (lochana):

if put values for x from negative to positive. you will notice that if x<-2 ; d(D)/d(x) is positive if x = -2 ;d(D)/d(x) is zero if -2 <x < 2 ;d(D)/d(x) is negative if x = 2 ;d(D)/d(x) is zero if x < 2 ;d(D)/d(x) positive

OpenStudy (lochana):

that's how I argue that fence has its minimum value when x = 2 and of course a distance cannot be an negative value anyway:). so we forget x = -2

OpenStudy (lochana):

now go back to our distance equation D = 2x + 8/x

OpenStudy (lochana):

put x = 2 D = 2*2 + 8/2 = 4 +4 = 8

OpenStudy (lochana):

hope this helps

OpenStudy (anonymous):

thanks!